Solve. Round results to the nearest thousandth. x2 + 8x - 33 = 0
Answer by stanbon(75887)

Solve. Round results to the nearest thousandth. x2 + 8x - 33 = 0
---
x^2 + 8x - 33 = 0
x^2 + 11x - 3x - 33 = 0
x(x+11)-3(x+11) = 0
(x+11)(x-3) = 0
x = -11 or x = 3
=====================
Cheers,
Stan H.